OpenStudy (idku):

So the formula for taylor series is ___, and how about 1/x^3?

OpenStudy (idku):

\[\color{blue}{y=f(x)~~~{\rm at~~x=a~~is~given~by~}~~\sum^{\infty}_{n=0}\frac{f^{(n)}(a)}{n!}(x-a)}\]

OpenStudy (fibonaccichick666):

so what's the how about for?

OpenStudy (idku):

I will try to do 1/x^3 at x=1

OpenStudy (fibonaccichick666):

ah, ok, so first, find some derivatives

OpenStudy (idku):

f(x)=1/x^3 f(1)=1 f'(x)=-3/x^4 f'(1)=-3 f''(x)=24/x^5 f''(1)=24 f'''(x)=-120/x^6 f'''(1)=-120

OpenStudy (fibonaccichick666):

how'd you get 24?

OpenStudy (fibonaccichick666):

check your second derivative

OpenStudy (idku):

-yeah, my bad it is 12, and then -60

OpenStudy (fibonaccichick666):

that twould be correct

OpenStudy (idku):

and what do I do after ?

OpenStudy (fibonaccichick666):

plug and chug

OpenStudy (fibonaccichick666):

a=1

OpenStudy (fibonaccichick666):

then there is a remainder theorem for the remaining

OpenStudy (fibonaccichick666):

but you just need to find the limit if I understand

OpenStudy (idku):

I think I am too weak for taylor series... not that I ever was good at math. I will review easier examples first and then do this one.

OpenStudy (idku):

tnx for your input

OpenStudy (fibonaccichick666):

wait, you've already done all the work

OpenStudy (rational):

\[\color{blue}{y=f(x)~~~{\rm at~~x=a~~is~given~by~}~~\sum^{\infty}_{n=0}\frac{f^{(n)}(a)}{n!}(x-a)^{\color{red}{n}}}\]

OpenStudy (fibonaccichick666):

\[ \sum^{\infty}_{n=0} \frac{f^{(n)}(a)}{n!}(x-a)=f(1)+f'(1)/1! *(x-1)+f"(1)/2!*(x-1)^2+...\ \]

OpenStudy (fibonaccichick666):

yea, forgot to toss in the exponent, sorry

OpenStudy (idku):

then I am pretty stupid that I can't plug it in, do I get? \[=1+\frac{(-3)}{1!}(x-1)+\frac{(12)}{2!}(x-1)^2+\frac{(-60)}{3!}(x-1)^3+\frac{(360)}{4!}(x-1)^4+...\]

OpenStudy (fibonaccichick666):

yea, that's all there is to it, just simplify

OpenStudy (idku):

\[+\frac{(360)}{4!}(x-1)^4+\frac{-360 \times -7}{5!}(x-1)^5+....\]

OpenStudy (fibonaccichick666):

then they like to use a remainder theorem

OpenStudy (fibonaccichick666):

usually they will say what power they want this in

OpenStudy (idku):

I haven't learned that yet. My teacher said the he doesn't consider it very important.

OpenStudy (idku):

Well, if it won't be a pain I can look it up and learn it....

OpenStudy (fibonaccichick666):

it's not crazy important, but you just did it

OpenStudy (idku):

I just did the remainder theory ?

OpenStudy (idku):

anyway... I will worry about it later. I need to comprehend the task of writing the taylor series (which is not as crazy as i though it was).

OpenStudy (idku):

tnx again. CLOSED

OpenStudy (fibonaccichick666):

np

OpenStudy (fibonaccichick666):

for your reference http://tutorial.math.lamar.edu/Classes/CalcII/TaylorSeries.aspx